if (k-3),(2k+1)and (4k+3) arte three consecutive terms of an a.p. find the value of k.

Grade:10

1 Answers

Arun
6 years ago
Dear Meenu
If the above three terms are in A.P.
Then
2k +1 -(k-3) = 4k+3 - (2k+1)
2 k +1 -k +3 = 4k +3 -2k -1
k + 4 = 2k +2
k = 2
Regards
Arun (askIITians forum expert)